OpenStudy (anonymous):

Jeremy wants to buy a new computer. The saleswoman says that he can make a down payment and then pay for the computer in installments. Here's a formula that describes this scenario: x=t-yz x = Amount down y = Money each month z = Number of months t = Total price Rewrite the formula to solve for the amount of money Jeremy must pay each month. @texaschic101 @Kristen17

OpenStudy (anonymous):

x + yz = t

OpenStudy (anonymous):

Sorry I didthe wrong one

OpenStudy (texaschic101):

your looking for y x = t - yz (subtract t from both sides) x - t = -yz (now divide by z) x/z - t/z = -y (multiply by -1) -x/z + t/z = y or (-x + t)/z = y

OpenStudy (anonymous):

y=t-x/z should be right

OpenStudy (anonymous):

\[\frac{ -x+t }{ z }=y\] Was correct, thanks @texaschic101 :) AGAIN! Haha

OpenStudy (texaschic101):

Kristen's answer is also correct...she just put the t first instead of the -x. It is still the same answer

OpenStudy (texaschic101):

glad to help :)